What is the domain and range to this function: 16y^2/9 + 4x^2/25 = 1
Answer by stanbon(75887)

What is the domain and range to this function:
16y^2/9 + 4x^2/25 = 1
---
Rewrite as:
y^2/(9/16) + x^2/(25/4) = 1
---------------
This is an ellipse with seim-major axis = (5/4)
and semi-minor axis = (3/4)
---
The center is at (0,0)
------------------------------
Domain: -5/4 <= x <= 5/4
Range: -3/4 <= y <= 3/4
==============================
Cheers,
Stan H.